[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jun 06, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 6 21:09:29 UTC 2016


HSU Shift Log for Mon Jun 06 2016 17:09:23 GMT-0400 (EDT)

Shift: H

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: UKMET tracker has not been in ATCF for Colin during the past  
couple of days.

N-AWIPS/AWIPS2 issues: None

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160606/05a7f7ab/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list